OpenStudy (anonymous):

the question is.. the perimeter of a rectangle equals the sum of the length of it's four sides, if the width is 5 and the perimeter is 36, what is the length

OpenStudy (anonymous):

 now perimeter=2(L+W) or L=(P/2)-W put the values in this equation.. can u do it?